Question No#13 Let h(π‘₯) = π‘₯2βˆ’2π‘₯βˆ’3

a) Make a table of the values of h(x) at x = 2.9, 2.99, 2.999, and so on. Then estimate limh(π‘₯). What estimate do you arrive at if you evaluate h at x = 3.1, 3.01, 3.001 ...

π‘₯β†’3

instead?

b) Support your conclusions in part (a) by graphing h near x = 3 and using that graph to

estimate h(x) on the graph as x approaching 3.

c) Find lim h(π‘₯) algebraically.

π‘₯β†’3



1
Expert's answer
2021-10-06T10:21:07-0400

a)

"f(2.9)=-0.39,\ f(2.99)=-0.0399,\ f(2.999)=-0.003999,"

"f(2.9999)=-0.00039999"

"\displaystyle{\lim_{x\to 3^-}}(x^2-2x-3)=0"


"f(3.1)=0.41,\ f(3.01)=0.0401,\ f(3.001)=0.004001,"

"f(3.0001)=0.00040001"

"\displaystyle{\lim_{x\to 3^+}}(x^2-2x-3)=0"


b)



c)

"\displaystyle{\lim_{x\to 3^-}}(x^2-2x-3)=3^2-2\cdot 3-3=0"